# Algorithmic Behaviors Across Regions: A Geolocation Audit of YouTube Search for COVID-19 Misinformation between the United States and South Africa

### v1

@@ Appendix I Significance Testing Details
Index Model Features Vectorizer Zero-/Few-Shot Metadata Temp. Acc. F1-M F1-W
0 SVM Title FastText - - - 0.64 0.64 0.64
1 SVM Description FastText - - - 0.6 0.6 0.6
2 SVM Transcript FastText - - - 0.58 0.58 0.58
3 SVM Tags FastText - - - 0.62 0.62 0.62
4 SVM Comments FastText - - - 0.49 0.47 0.47
5 SVM Title,Description FastText - - - 0.68 0.67 0.68
6 SVM Title,Transcript FastText - - - 0.71 0.71 0.71
7 SVM Title,Tags FastText - - - 0.71 0.71 0.71
8 SVM Title,Comments FastText - - - 0.66 0.66 0.66
9 SVM Title,Description,Transcript FastText - - - 0.69 0.69 0.69
10 SVM Title,Description,Tags FastText - - - 0.68 0.68 0.68
11 SVM Title,Description,Comments FastText - - - 0.66 0.66 0.66
12 SVM Title,Description,Transcript,Tags FastText - - - 0.69 0.69 0.69
13 SVM Title,Description,Transcript,Comments FastText - - - 0.67 0.67 0.67
14 SVM Title,Description,Transcript,Tags,Comments Count(unigram) - - - 0.64 0.64 0.64
15 SVM Title,Description,Transcript,Tags,Comments Count(unigram,bigram) - - - 0.68 0.68 0.68
16 SVM Title,Description,Transcript,Tags,Comments Count(unigram,bigram,trigram) - - - 0.68 0.69 0.69
17 SVM Title,Description,Transcript,Tags,Comments TFIDF(unigram) - - - 0.78 0.78 0.78
18 SVM Title,Description,Transcript,Tags,Comments TFIDF(unigram,bigram) - - - 0.78 0.78 0.78
19 SVM Title,Description,Transcript,Tags,Comments TFIDF(unigram,bigram,trigram) - - - 0.77 0.76 0.77
20 SVM Title,Description,Transcript,Tags,Comments Word2Vec - - - 0.76 0.76 0.76
21 SVM Title,Description,Transcript,Tags,Comments FastText - - - 0.71 0.71 0.72
22 SVM Title,Description TFIDF(unigram) - - - 0.7 0.7 0.7
23 SVM Title,Description TFIDF(unigram,bigram) - - - 0.72 0.72 0.72
24 SVM Title,Transcript TFIDF(unigram) - - - 0.73 0.73 0.73
25 SVM Title,Transcript TFIDF(unigram,bigram) - - - 0.75 0.75 0.75
26 SVM Title,Tags TFIDF(unigram) - - - 0.7 0.7 0.7
27 SVM Title,Tags TFIDF(unigram,bigram) - - - 0.71 0.71 0.71
28 SVM Title,Description,Transcript TFIDF(unigram) - - - 0.73 0.73 0.73
29 SVM Title,Description,Transcript TFIDF(unigram,bigram) - - - 0.76 0.76 0.76
30 SVM Title,Description,Tags TFIDF(unigram) - - - 0.72 0.72 0.72
31 SVM Title,Description,Tags TFIDF(unigram,bigram) - - - 0.75 0.75 0.76
32 SVM Title,Description,Transcript,Tags TFIDF(unigram) - - - 0.75 0.75 0.75
33 SVM Title,Description,Transcript,Tags TFIDF(unigram,bigram) - - - 0.78 0.78 0.78
34 SVM Title,Description,Transcript,Comments TFIDF(unigram) - - - 0.75 0.75 0.75
35 SVM Title,Description,Transcript,Comments TFIDF(unigram,bigram) - - - 0.76 0.76 0.76
36 XGB Title,Description TFIDF(unigram) - - - 0.68 0.68 0.68
37 XGB Title,Description TFIDF(unigram,bigram) - - - 0.67 0.67 0.67
38 XGB Title,Transcript TFIDF(unigram) - - - 0.71 0.71 0.71
39 XGB Title,Transcript TFIDF(unigram,bigram) - - - 0.69 0.69 0.69
41 XGB Title,Tags TFIDF(unigram) - - - 0.69 0.69 0.69
42 XGB Title,Tags TFIDF(unigram,bigram) - - - 0.69 0.69 0.69
43 XGB Title,Description,Transcript TFIDF(unigram) - - - 0.7 0.69 0.69
44 XGB Title,Description,Transcript TFIDF(unigram,bigram) - - - 0.73 0.73 0.73
45 XGB Title,Description,Tags TFIDF(unigram) - - - 0.74 0.74 0.74
46 XGB Title,Description,Tags TFIDF(unigram,bigram) - - - 0.75 0.75 0.76
47 XGB Title,Description,Transcript,Tags TFIDF(unigram) - - - 0.71 0.71 0.71
48 XGB Title,Description,Transcript,Tags TFIDF(unigram,bigram) - - - 0.72 0.72 0.72
49 XGB Title,Description,Transcript,Comments TFIDF(unigram) - - - 0.73 0.73 0.73
50 XGB Title,Description,Transcript,Comments TFIDF(unigram,bigram) - - - 0.72 0.72 0.72
51 XGB Title,Description,Transcript,Tags,Comments TFIDF(unigram) - - - 0.73 0.73 0.73
52 XGB Title,Description,Transcript,Tags,Comments TFIDF(unigram,bigram) - - - 0.73 0.73 0.73
53 DeBerta (base) Title,Description,Transcript,Tags - - - - 0.81 0.81 0.81
54 DeBerta (large) Title,Description,Transcript,Tags - - - - 0.85 0.85 0.85
@@
55 GPT-4 Turbo Title,Description,Transcript,Tags - Zero-Shot Complete 0 0.76 0.76 0.76
@@
56 GPT-4 Turbo Title,Description,Transcript,Tags - Zero-Shot Truncated 0 0.78 0.78 0.78
@@
57 GPT-4 Turbo Title,Description,Transcript,Tags - Zero-Shot Truncated 0.2 0.79 0.79 0.79
@@
58 GPT-4 Turbo Title,Description,Transcript,Tags - Zero-Shot Truncated 0.7 0.77 0.77 0.77
59 GPT-4 Turbo Title,Description,Transcript,Tags - Few-Shot Truncated 0 0.78 0.78 0.78
60 GPT-4 Turbo Title,Description,Transcript,Tags - Few-Shot Truncated 0.2 0.78 0.78 0.78
61 GPT-4 Turbo Title,Description,Transcript,Tags - Few-Shot Truncated 0.7 0.79 0.79 0.79
@@
Table 10: Performances on our 10% held-out test set (273 videos) across Support Vector Machines (SVM), XGBoost (XGB), DeBerta-v3-base, DeBerta-v3-large, and GPT-4 Turbo (Version 1106) on various features, vectorizers, and parameters. We trained and evaluated 62 different models. For GPT-4 prompts, we experimented with providing complete vs. truncated video metadata (first 500 tokens of each video metadata) in the prompts; these are indicated by the values “Truncated” and “Complete” within the column “Metadata.” The highest performance value is bolded; DeBerta-v3-large (Index 54) performed the best, obtaining 0.85 accuracy, weighted F1 score, and macro F1 score. Note that Temp: Temperature, Acc.: Accuracy, F1-M: Macro F1-score, F1-W: Weighted F1-score.
@@
Table 11: Zero-shot prompt to determine whether a YouTube video opposes COVID-19 misinformation, supports COVID-19 misinformation, or is neutral information. For each label, we employed the same descriptions provided to AMT workers on their annotation task. Note that “Neutral Information” here refers to the merged category, consisting of “Neutral COVID-19 Information,” “COVID-19 origins,” and “Irrelvant” labels. This merged category includes information that neither supports nor opposes COVID-19 misinformation. For more information, please see subsection “Consolidating from 5-classes to 3-classes” in §5.
@@
Table 12: Few-shot prompt to determine whether a YouTube video opposes COVID-19 misinformation, supports COVID-19 misinformation, or is neutral information. For each label, we employed the same descriptions and examples provided to AMT workers on their annotation task. Note that “Neutral Information” here refers to the merged category, consisting of “Neutral COVID-19 Information,” “COVID-19 origins,” and “Irrelvant” labelsNote that “Neutral Information” here refers to the merged category, consisting of “Neutral COVID-19 Information,” “COVID-19 origins,” and “Irrelvant” labels. This merged category includes information that neither supports nor opposes COVID-19 misinformation. For more information, please see subsection “Consolidating from 5-classes to 3-classes” in §5.

### v2

@@ Appendix I Significance Testing Details
Condition p r Mean Rank Diff.
Main Finding *** 0.49 SA $>$ US
Case 1 *** 0.57 SA $>$ US
Case 2 ** 0.37 SA $>$ US
Table 9: Summary of the main finding and validation checks under two alternative treatments of the “On the COVID-19 origins in Wuhan, China” label. The results are based on Mann-Whitney U Tests comparing misinformation bias scores between the US and SA for the top-10 videos in the SERPs. Each test reports the p-value (p), Mann-Whitney effect size (r), and mean rank difference. For example, SA $>$ US indicates that bots in SA received more misinformative SERPs than bots in the US. These validation checks reinforce the robustness of our main findings, showing consistent results regardless of how lab leak theory videos were categorized. Note that: *$P<0.05$; **$P<0.01$; ***$P<0.001$.
Robustness Analysis Under Alternative Treatments of Lab Leak Theory Class
Since the Lab Leak Theory are often featured heavily as misinformation, we conducted additional validation checks to evaluate the robustness of our main findings under two alternative treatments of the “On the COVID-19 origins in Wuhan, China” label. Originally, videos discussing COVID-19 origins were labeled as “On the COVID-19 origins in Wuhan, China”(see §4) and later merged into a neutral “neither” class to reflect the lack of consensus about their status as misinformation. [^21] As of September 2024, Abdelfatah and Arablouei (2024) maintained that the Lab Leak Theory is plausible. Considering the top-10 videos in the SERPs, our main finding showed that bots in SA were exposed to significantly more misinformative SERPs than bots in the US (p $<$ 0.001, effect size r=0.49). We replicate this main finding by conducting the same significance test while 1) excluding videos labeled under the “On the COVID-19 origins in Wuhan, China” class in the SERPs (Case 1), and 2) treating the class as misinformation (Case 2). Table 9 summarizes the results of the main finding and validation checks, demonstrating consistent results across all scenarios, regardless of how lab leak theory videos were categorized.
Case 1: Eliminating the class. We removed all videos labeled as “On the COVID-19 origins in Wuhan, China” from the dataset and repeated the significance test to compare the misinformation bias scores for the top-10 search results between the US and SA. Our results confirmed that bots in SA still encountered significantly more misinformative SERPs compared to bots in the US (p $<$ 0.001, effect size r=0.57), consistent with our main findings.
Case 2: Merging the class with misinformation. We merged the “On the COVID-19 origins in Wuhan, China” class with the “supporting misinformation” class and conducted the same significance test. Again, SA bots were exposed to significantly more misinformative SERPs compared to bots in the US (p $<$ 0.01, effect size r=0.37), consistent with our main findings.
These validation checks reinforce the robustness of our findings, demonstrating that bots in SA encountered significantly more misinformative SERPs than bots in the US regardless of how lab leak theory videos were categorized.
Scale Annotation Label Descriptions Example Video (Title and ID)
-1 Opposing COVID-19 Misinformation (Opposing) The video opposes, debunks, satirizes OR provides countervailing contexts to COVID-19 misinformation OR disseminates/endorses public health authorities’ policies OR explains the scientific contexts surrounding the public health policy and the COVID-19 pandemic OR provides countervailing contexts to the usage of misleading terms, such as “Kungflu” and “China Virus”. 1. Fox Allows Lunatic to Imply Kim Jong-Un Created Cronavirus (Video ID: 8EGVzWRic1I) 2. The Monday Times: Coronavirus Car, Ghosts, Tik Tok, 5G Conspiracy — The Daily Show With Trevor Noah (Video ID: 2totKKH4770)
0 Neutral COVID-19 Information (Neutral) Videos that broadly cover anything related to the COVID-19 pandemic, but do not support nor oppose COVID-19 misinformation. For example, videos that report on the number of COVID-19 cases, discuss remote work hours during a pandemic, etc. 1. Coronavirus outbreak: Hospitals in China swamped with patients (Video ID: 9qAFanPeKtE) 2. Africa Unite Against COVID 19 (Video ID: 6EMEEn65JPg)
1 Supporting COVID-19 Misinformation (Supporting) The video supports OR provides evidence for COVID-19 misinformation OR suggests alternative treatment, prevention, and cure without scientific evidence OR exaggerates a claim to suggest a misinformative narrative OR promotes the usage of misleading terms without countervailing contexts. 1. KT The Arch Degree Break’s Down The Coronavirus! (MUST SEE) (Video ID: ulncPbTgPlo) 2. Origin of the CCP Virus: Fauci Emails Revealed; Ivermectin: A Vaccine Alternative?(Video ID: edT1HYMjnzA)
2 On the COVID-19 origins in Wuhan, China (COVID-19 origins) Any videos that hypothesize OR cover the origins of COVID-19 in Wuhan, China as the main premise of the video (e.g. natural origins theory, lab leak theory) OR provide evidence for the lab leak theory OR trying to debunk/satirize the lab leak theory. Videos that speculate beyond the origin theories of the virus (e.g., COVID-19 was intentionally engineered as a bioweapon in the Wuhan Lab) are marked as supporting COVID-19 misinformation. 1. Progressive Lab Leak Theory (Video ID: 0iNZE6BEoXs) 2. Why the US is taking a second look at the ’lab leak’ theory about COVID-19 (Video ID: ZRV_uYS2Buc)
3 Irrelevant Any video whose content is not related to the COVID-19 pandemic and vaccines. 1. Dengue: Laboratory-acquired case reported in North Carolina, according to study (Video ID: PxFJ3YSgG0Y) 2. The Nuremberg Code (Video ID: Eu_3Cbf7f8Y)
4 YouTube video in a language other than English (Non-English) The video’s title, description, and/or contents in a language other than English. 1. Can Coronavirus be Biological Weapon? Video Analysis by Major General S B Asthana (Retd) (Video ID: WiHmN_DuQhQ) 2. Shiva $\mid$ The Arrogant Kung Fu Fighter $\mid$ Episode 54 $\mid$ Download Voot Kids App (Video ID: ISfm5pAWmxs)
ID Model Features Vectorizer Zero-/Few-Shot Metadata Temp. Acc. F1-M F1-W
SVM with Individual Features (FastText)
0 SVM Title FastText 0.64 0.64 0.64
1 SVM Description FastText 0.60 0.60 0.60
2 SVM Transcript FastText 0.58 0.58 0.58
3 SVM Tags FastText 0.62 0.62 0.62
4 SVM Comments FastText 0.49 0.47 0.47
SVM with Feature Combinations (FastText)
5 SVM Title, Description FastText 0.68 0.67 0.68
6 SVM Title, Transcript FastText 0.71 0.71 0.71
7 SVM Title, Tags FastText 0.71 0.71 0.71
8 SVM Title, Comments FastText 0.66 0.66 0.66
9 SVM Title, Description, Transcript FastText 0.69 0.69 0.69
10 SVM Title, Description, Tags FastText 0.68 0.68 0.68
11 SVM Title, Description, Comments FastText 0.66 0.66 0.66
12 SVM Title, Description, Transcript, Tags FastText 0.69 0.69 0.69
13 SVM Title, Description, Transcript, Comments FastText 0.67 0.67 0.67
SVM with All Features (Various Vectorizers)
14 SVM All Features Count(unigram) 0.64 0.64 0.64
15 SVM All Features Count(unigram, bigram) 0.68 0.68 0.68
16 SVM All Features Count(unigram, bigram, trigram) 0.68 0.69 0.69
17 SVM All Features TFIDF(unigram) 0.78 0.78 0.78
18 SVM All Features TFIDF(unigram, bigram) 0.78 0.78 0.78
19 SVM All Features TFIDF(unigram, bigram, trigram) 0.77 0.76 0.77
20 SVM All Features Word2Vec 0.76 0.76 0.76
21 SVM All Features FastText 0.71 0.71 0.72
SVM with Feature Combinations (TFIDF)
22 SVM Title, Description TFIDF(unigram) 0.70 0.70 0.70
23 SVM Title, Description TFIDF(unigram, bigram) 0.72 0.72 0.72
24 SVM Title, Transcript TFIDF(unigram) 0.73 0.73 0.73
25 SVM Title, Transcript TFIDF(unigram, bigram) 0.75 0.75 0.75
26 SVM Title, Tags TFIDF(unigram) 0.70 0.70 0.70
27 SVM Title, Tags TFIDF(unigram, bigram) 0.71 0.71 0.71
28 SVM Title, Description, Transcript TFIDF(unigram) 0.73 0.73 0.73
29 SVM Title, Description, Transcript TFIDF(unigram, bigram) 0.76 0.76 0.76
30 SVM Title, Description, Tags TFIDF(unigram) 0.72 0.72 0.72
31 SVM Title, Description, Tags TFIDF(unigram, bigram) 0.75 0.75 0.76
32 SVM Title, Description, Transcript, Tags TFIDF(unigram) 0.75 0.75 0.75
33 SVM Title, Description, Transcript, Tags TFIDF(unigram, bigram) 0.78 0.78 0.78
34 SVM Title, Description, Transcript, Comments TFIDF(unigram) 0.75 0.75 0.75
35 SVM Title, Description, Transcript, Comments TFIDF(unigram, bigram) 0.76 0.76 0.76
XGBoost with Feature Combination (TFIDF)
36 XGB Title, Description TFIDF(unigram) 0.68 0.68 0.68
37 XGB Title, Description TFIDF(unigram, bigram) 0.67 0.67 0.67
38 XGB Title, Transcript TFIDF(unigram) 0.71 0.71 0.71
39 XGB Title, Transcript TFIDF(unigram, bigram) 0.69 0.69 0.69
41 XGB Title, Tags TFIDF(unigram) 0.69 0.69 0.69
42 XGB Title, Tags TFIDF(unigram, bigram) 0.69 0.69 0.69
43 XGB Title, Description, Transcript TFIDF(unigram) 0.70 0.69 0.69
44 XGB Title, Description, Transcript TFIDF(unigram, bigram) 0.73 0.73 0.73
45 XGB Title, Description, Tags TFIDF(unigram) 0.74 0.74 0.74
46 XGB Title, Description, Tags TFIDF(unigram, bigram) 0.75 0.75 0.76
47 XGB Title, Description, Transcript, Tags TFIDF(unigram) 0.71 0.71 0.71
48 XGB Title, Description, Transcript, Tags TFIDF(unigram, bigram) 0.72 0.72 0.72
49 XGB Title, Description, Transcript, Comments TFIDF(unigram) 0.73 0.73 0.73
50 XGB Title, Description, Transcript, Comments TFIDF(unigram, bigram) 0.72 0.72 0.72
51 XGB All Features TFIDF(unigram) 0.73 0.73 0.73
52 XGB All Features TFIDF(unigram, bigram) 0.73 0.73 0.73
DeBERTa Models
53 DeBERTa (base) Title, Description, Transcript, Tags 0.81 0.81 0.81
54 DeBERTa (large) Title, Description, Transcript, Tags 0.85 0.85 0.85
GPT-4 Turbo Models
@@
55 GPT-4 Turbo Title, Description, Transcript, Tags Zero-Shot Complete 0 0.76 0.76 0.76
@@
56 GPT-4 Turbo Title, Description, Transcript, Tags Zero-Shot Truncated 0 0.78 0.78 0.78
@@
57 GPT-4 Turbo Title, Description, Transcript, Tags Zero-Shot Truncated 0.2 0.79 0.79 0.79
@@
58 GPT-4 Turbo Title, Description, Transcript, Tags Zero-Shot Truncated 0.7 0.77 0.77 0.77
59 GPT-4 Turbo Title, Description, Transcript, Tags Few-Shot Truncated 0 0.78 0.78 0.78
60 GPT-4 Turbo Title, Description, Transcript, Tags Few-Shot Truncated 0.2 0.78 0.78 0.78
61 GPT-4 Turbo Title, Description, Transcript, Tags Few-Shot Truncated 0.7 0.79 0.79 0.79
@@
Table 11: Performances on our held-out test set (273 videos) across Support Vector Machines (SVM), XGBoost (XGB), DeBERTa-v3, and GPT-4 Turbo (Version 1106) on various features, vectorizers, and parameters. We trained and evaluated 62 models. For GPT-4, we experimented with providing complete vs. truncated video metadata (first 500 tokens of each video metadata) in the prompts; these are indicated by the values “Truncated” and “Complete” in the column “Metadata.” The highest performance value is bolded; DeBERTa-v3-large (Index 54) performed the best. Note that Temp: Temperature, Acc.: Accuracy, F1-M: Macro F1-score, F1-W: Weighted F1-score, and All Features: Title, Description, Transcript, Tags, Comments.
@@
Table 12: Zero-shot prompt to determine whether a YouTube video opposes COVID-19 misinformation, supports COVID-19 misinformation, or is neutral information. For each label, we employed the same descriptions provided to AMT workers on their annotation task. Note that “Neutral Information” here refers to the merged category, consisting of “Neutral COVID-19 Information,” “COVID-19 origins,” and “Irrelvant” labels. Please see subsection “Consolidating from 5-classes to 3-classes” in §5 for more information.
@@
Table 13: Few-shot prompt to determine whether a YouTube video opposes COVID-19 misinformation, supports COVID-19 misinformation, or is neutral information. For each label, we employed the same descriptions and examples provided to AMT workers on their annotation task. Note that “Neutral Information” here refers to the merged category, consisting of “Neutral COVID-19 Information,” “COVID-19 origins,” and “Irrelvant” labels. Please see subsection “Consolidating from 5-classes to 3-classes” in §5 for more information.